If a consumer is consuming two commodity X and Y and his utility function U(X,Y)=2xy+6.if the price of the two commodity are 2 and 6 respectively, and consumer has a total income of 80 birr to be spent on the two goods,a)find the utility maximizing Quantity of good X and Y (b) find the MRSxy at equilibrium (c) find the MRSyx at equilibrium

1
Expert's answer
2022-02-06T14:22:51-0500

Solution:

a.). Utility maximizing quantity = MRSXY = "\frac{MUx}{MUy} = \frac{Px}{Py}"

MUx = "\frac{\partial U} {\partial X}" = 2y


MUy = "\frac{\partial U} {\partial Y}" = 2x

Px = 2

Py = 6

"\frac{MUx}{MUy} = \frac{Px}{Py}"


"\frac{2y}{2x} = \frac{2}{6}"

Y = "\frac{x}{3}"

Substitute the figure of X in the budget constraint:

Budget constraint: M = PxX + PyY

80 = 2x + 6y

80 = 2x + 6("\frac{x}{3}") = 2x + 2x = 4x

80 = 4x

x = 20

y = "\frac{x}{3}" = "\frac{20}{3}" = 6.67

The utility maximizing quantity of X and Y = (20, 6.67)

 

b.). MRSxy at equilibrium:

MRSxy = "\frac{MUx}{MUy} = \frac{Px}{Py}"


="\frac{MUx}{Px} = \frac{MUy}{Py}"

X = 3y

 

c.). MRSyx at equilibrium:

MRSyx ="\frac{MUy}{MUx} = \frac{Py}{Px}"


= "\frac{2x}{6} = \frac{2y}{2}"


Y = "\frac{x}{3}"